### v1

# fastrerandomize: An R Package for Fast Rerandomization Using Accelerated Computing

Abstract
The fastrerandomize R package provides hardware-accelerated tools for performing rerandomization and randomization testing in experimental research. Using a JAX backend, the package enables exact rerandomization inference even for large experiments with hundreds of billions of possible randomizations. Key functionalities include generating pools of acceptable rerandomizations based on covariate balance, conducting exact randomization tests, and performing pre-analysis evaluations to determine optimal rerandomization acceptance thresholds. Through batched processing and GPU acceleration, fastrerandomize achieves substantial performance gains compared to existing implementations, making previously intractable designs computationally feasible. The package therefore extends the randomization-based inference toolkit in R, allowing researchers to efficiently implement more stringent rerandomization designs and conduct valid inference even with large sample sizes or in high-dimensional settings.
keywords:
Rerandomization, randomization tests, experimental design, hardware acceleration, R
Introduction
Rerandomization is a powerful strategy for ensuring strong covariate balance in experimental research. By repeatedly drawing treatment assignments until a prespecified balance criterion is met, rerandomization helps improve the precision of estimated treatment effects relative to pure random assignment (Morgan and Rubin, 2012; Li et al., 2018). Despite its advantages, rerandomization poses a formidable computational challenge: as the size of an experiment grows, the space of possible assignments can quickly reach hundreds of billions, complicating both the generation of acceptable randomizations and the conduct of valid inference. Although theoretical work has established the statistical validity of rerandomization, user-friendly software that efficiently implements these methods in large-scale settings has, until now, been more limited.
The fastrerandomize package addresses this need by providing hardware-accelerated tools for rerandomization and randomization testing within R. By leveraging JAX, fastrerandomize supports hardware backends such as CPU, CUDA, and METAL, enabling researchers to conduct exact randomization-based inference even at large experimental scales. Key functionalities include (1) generating pools of acceptable assignments based on desired balance thresholds, (2) running exact randomization tests with the accepted pools, and (3) performing pre-analysis evaluations to guide the choice of optimal rerandomization thresholds.
In what follows, we provide an overview of the statistical framework underpinning fastrerandomize, detail its core functionalities, and illustrate how it may be applied in practice. After literature review in Section 1, Section 2 outlines the primary methodologies implemented in the package, Section 3 describes its software architecture and usage. Sections 4 and 5 offer examples demonstrating both basic and more advanced analyses. We conclude with potential directions for future development in Section 7.
Figure 1: Core workflow of the fastrerandomize package for conducting rerandomization and randomization testing. The process begins with covariate data and proceeds through five main steps: (1) generating candidate randomizations using compiled code, (2) checking covariate balance with GPU-accelerated computations, (3) accepting or rejecting assignments using Monte Carlo methods for large experiments, (4) conducting exact randomization tests using the pool of accepted randomizations, and (5) optionally computing fiducial intervals. Steps are implemented with hardware acceleration to enable efficient processing of large-scale experiments.
1 Literature Review and Related Packages
Over the past decade, an expanding body of open-source software has emerged to facilitate experimental design, with a particular focus on random assignment procedures and randomization-based inference. Many of these tools incorporate rerandomization techniques to improve covariate balance across treatment groups, thereby enhancing the statistical power of a study (Morgan and Rubin, 2012; Li et al., 2018). Despite the proliferation of solutions, existing packages often address only narrow classes of experiments and frequently encounter computational bottlenecks in large-scale settings with many units or covariates. This section surveys some of the most widely used packages for general random assignment, randomization testing, and rerandomization, before discussing the core challenges that persist in these domains.
1.1 General Random Assignment and Testing Packages
A variety of open-source packages are available for general-purpose random assignment and randomization-based testing. The randomizr package, for instance, streamlines the assignment of study units to diverse treatment groups (Coppock, 2024). Another solution, RATest, incorporates randomization tests applicable to multiple experimental design scenarios and offers sample datasets that illustrate, for example, permutation testing in regression discontinuity contexts (Olivares-Gonzalez and Sarmiento-Barbieri, 2017).
The RCT2 package assists with planning and analyzing two-stage randomized controlled trials, providing hypothesis tests for direct and spillover effects alongside tools for power calculations (Huang et al., 2022). Additional general-purpose frameworks include randomizeR (Coppock, 2024), which addresses both the planning and analysis phases of randomized experiments, as well as RItools and ri2 (Coppock, 2022; Bowers et al., 2023), which support randomization inference under a range of assignment schemes (including complete, block, and cluster randomization).
1.2 Rerandomization and Testing Packages
In a similar vein, rerandomization has attracted growing attention because of its ability to improve covariate balance and thus increase the precision of estimated treatment effects (Morgan and Rubin, 2012; Branson et al., 2024; Kapelner et al., 2022). A smaller set of available software specifically focuses on rerandomization. The rerandPower package offers power analyses and sample size calculations to inform rerandomized designs (Branson et al., 2024).
Meanwhile, OptimalRerandExpDesigns helps users identify rerandomization thresholds that account for the distribution of covariates and user-defined balance criteria (Kapelner et al., 2022). The jumble package supports iterative generation of random assignments for experiments with up to two treatment arms, based on user-specified acceptance probabilities and balance measures (McConeghy, 2024). Finally, rct facilitates balanced random assignments for randomized controlled trials with any number of treatment arms, which makes it appealing for multifaceted experiments where binary-treatment approaches may be unsuitable (Banerjee et al., 2017).
1.3 Challenges and Gaps
Although these packages advance the design and analysis of experiments, several limitations constrain their application to large-scale studies. One persistent concern is that many solutions can verify or test an existing randomization but provide limited functions for the initial task of constructing assignment schemes automatically. This requires researchers to generate and manage their own candidate pools of randomizations.
Scalability also presents a bottleneck in rerandomization: as experiments grow to include more units and covariates, generating, evaluating, and storing large pools of acceptable randomizations becomes computationally intensive. Current software packages often struggle when handling scenarios where investigators seek stringent acceptance thresholds with hundreds of billions of potential randomizations, or in experiments involving hundreds or thousands of covariates—a scenario becoming increasingly common with the rise of causal machine learning involving high-dimensional text (Keith et al., 2020), image (Jerzak et al., 2023), or network data (Ogburn et al., 2024).
While modern hardware offers parallelism gains through accelerated computing, existing implementations have not leveraged these advances. The fastrerandomize package addresses these limitations by automating balanced assignment generation and harnessing this acceleration, thereby making rerandomization and randomization-based inference both computationally efficient and practically feasible for large-scale experimental research.
2 Statistical Methods
The core statistical methodology in fastrerandomize adapts the rerandomization framework of Morgan and Rubin (2012), which repeatedly draws treatment assignments until achieving sufficient covariate balance. Let $W$ represent a vector of treatment assignments, $X$ a matrix of pre-treatment covariates, and $Y$ observed outcomes. A rerandomization design proceeds by generating candidate randomizations $W^{(1)},\ldots,W^{(M)}$, accepting each $W^{(m)}$ if meeting a balance criterion (e.g., Hotelling’s $T^{2}$) associated with an acceptance probability $p_{a}(X,W^{(m)})$. As illustrated in Figure 2, each candidate assignment is evaluated against a pre-specified balance criterion that yields an acceptance probability $p_{a}$. The smaller the acceptance probability, the more stringent the balance required before accepting a randomization.
Figure 2: Overview of the randomization acceptance process in rerandomization designs. Each candidate assignment is assessed based on covariate balance (e.g., Hotelling’s $T^{2}$), and rejected if not meeting a pre-selected threshold. This process continues until the targeted pool of acceptable randomizations is reached. (a) shows how with rerandomization, covariate balance improves. (b) shows that a stringent rejection threshold leads to low acceptance probability, necessitating more randomization draws.
Theory ensures that randomization inference remains valid as long as each accepted assignment is included with its correct probability in the final randomization distribution (Morgan and Rubin, 2012; Li et al., 2018). Because each accepted assignment is included with the correct probability (conditioned on acceptance), p-values and confidence intervals derived via randomization tests preserve nominal Type I error rates and correct coverage probabilities.
Once the pool of accepted randomizations is generated, a randomization test can be conducted by calculating a test statistic $T(W^{\textrm{Obs}},Y^{\textrm{Obs}})$ for the observed data and comparing it to the distribution of $T(W^{(m)},Y^{\textrm{Obs}})$ over the accepted pool, thereby yielding an exact p-value that respects rerandomization.
More formally, let $\mathcal{W}_{\text{Acc}}=\{W^{(1)},\dots,W^{(M_{\textrm{Acc}})}\}$ denote the final set of accepted randomizations, each meeting the acceptance threshold under the rerandomized design. Suppose we have a test statistic
$T(W,Y).$
The observed value actually obtained in the experiment is:
$T\bigl{(}W^{\mathrm{Obs}},Y^{\mathrm{Obs}}\bigr{)}.$
An exact two-sided p-value for the sharp null of no treatment effects (i.e., each unit’s potential outcomes are identical across treatment conditions) is given by
[…]
where $\mathbf{1}\{\cdot\}$ is the indicator function, and $M_{\textrm{Acc}}=|\mathcal{W}_{\text{Acc}}|$ is the number of accepted randomizations. Because $\mathcal{W}_{\text{Acc}}$ has been constructed to reflect the correct sampling probabilities under rerandomization, $p$ is a valid p-value that respects the fact that only certain assignments were accepted. In randomization-based inference, $p$-values in rerandomization tests are lower bounded by 1# of accepted $\frac{1}{\text{\# of accepted randomizations}}$ (Lehmann and Romano, 2005). If the acceptance threshold gets less and less stringent, this design-based p-value converges to what would be obtained under complete randomization (Morgan and Rubin, 2012; Li et al., 2018). Rerandomization also supports the construction of fiducial intervals by inversion of the test (Leemis, 2020).
The acceptance threshold for covariate balance has important implications for both statistical and computational performance. Tighter thresholds—leading to very low acceptance probabilities—can substantially improve covariate balance and thereby increase the efficiency (i.e., smaller variance) of estimated treatment effects. However, in practical terms, extremely stringent thresholds demand large candidate pools (and thus heavier computation) to obtain a sufficiently large set of accepted randomizations for stable inference. This challenge motivates the core contribution of this package.
Conversely, overly lenient thresholds can diminish the advantages of rerandomization by accepting many poorly balanced designs, causing the method to converge back to classical complete randomization. Accordingly, researchers often iterate over plausible acceptance probabilities in “pre-analysis” simulations to gauge how different thresholds affect downstream power, $p$-values, or interval widths (Branson et al., 2024; Kapelner et al., 2022). Such simulations also help ensure that the final pool of accepted randomizations is large enough for reliable test-based inferences and for constructing credible fiducial intervals, particularly in high-dimensional covariate settings (e.g., text, image, or network data (Keith et al., 2020; Jerzak et al., 2023; Ogburn et al., 2024)). [^1] In high-dimensional applications, one can improve numerical stability by incorporating penalized or regularized covariance estimates (e.g., diagonal or ridge-regularized inverses) into the balance statistic, ensuring the acceptance criterion continues to reflect meaningful imbalances.
3 Computational Methods
Implementing rerandomization at scale involves navigating an inherently combinatorial process, in which the total number of possible treatment assignments can be enormous. To address these computational challenges, fastrerandomize employs a hybrid of exact enumeration (for smaller designs) and Monte Carlo sampling (for large designs) while leveraging batched processing for memory efficiency. Central to this approach is the use of key-based storage, which replaces the explicit storage of every randomization vector with small keys that allow the randomizations to be regenerated on demand. Additionally, approximate techniques (e.g., diagonal covariance approximations) expedite balance checking in high-dimensional covariate spaces. Finally, the package exploits just-in-time (JIT) compilation and hardware acceleration via GPUs or specialized backends to further reduce run times. The methods outlined here enable fastrerandomize to handle rerandomization tasks that would be computationally prohibitive using naive approaches, even when the number of units, covariates, or candidate randomizations is very large.
3.1 Exact Enumeration and Batched Processing
The fastrerandomize package addresses a central challenge in rerandomization: balancing exactness with practical computational demands. To tackle this challenge, the package provides two primary strategies for generating and storing randomizations—exact enumeration and Monte Carlo sampling—each designed to accommodate different experimental sizes and memory constraints. Both approaches are implemented in JAX and exploit a batched workflow to reduce memory overhead, while R-facing functions manage data validation, type conversion, and user interaction. This integrated approach enables researchers to efficiently construct pools of acceptable randomizations across a range of experimental scales.
Exact Enumeration.
For smaller studies where the total number of possible assignments, $\binom{n}{n_{\textrm{Treated}}},$ is sufficiently small (e.g., up to a few hundred million), it is possible to exhaustively enumerate every distinct randomization. Users can invoke exact enumeration by specifying:
generate_randomizations(
n_units = 10,
n_treated = 5,
X = matrix(rnorm(10*3), nrow=10),
randomization_accept_prob = 0.1,
randomization_type = "exact"
This approach systematically iterates over all assignments, providing the maximum level of precision and guaranteeing that the entire randomization space is captured. Although exact enumeration is attractive from a statistical perspective, it becomes infeasible for even moderate sample sizes (e.g., $n=50$), because the number of possible assignments grows combinatorially.
Monte Carlo Sampling.
For larger experiments, fastrerandomize supports a Monte Carlo mechanism that draws candidate assignments in batches and evaluates their covariate balance:
big_covariate_matrix <- matrix(rnorm(500*100), nrow=500)
generate_randomizations(
n_units = nrow(big_covariate_matrix),
n_treated = round(nrow(big_covariate_matrix)/2),
X = big_covariate_matrix,
randomization_type = "monte_carlo",
max_draws = 5e5,
randomization_accept_prob = 0.001
Under Monte Carlo sampling, candidate assignments are randomly generated and checked against a user-specified criterion (e.g., Hotelling’s $T^{2}$). Those randomizations meeting the threshold are accepted. This approach can rapidly explore massive assignment spaces without enumerating them in full and is typically favored for large $n$.
Batched Processing and Memory Management.
Regardless of whether the randomizations are generated by exact enumeration or Monte Carlo sampling, fastrerandomize relies on batching to handle large-scale experiments efficiently:
  • Chunked Generation. Both the exhaustive listing of assignments and the Monte Carlo draws can be subdivided into user-defined batches (controlled by batch_size). This prevents the software from loading all candidate randomizations at once into memory.
  • Flexible Return Options. Users can further reduce memory loads by writing accepted assignments directly to disk rather than returning them to R (using file in generate_randomizations()). This is especially helpful when generating hundreds of thousands or millions of accepted randomizations.
By selecting exact enumeration for small designs and Monte Carlo with batching for large ones, analysts can scale rerandomization from modest experiments to those involving thousands of units. This unified approach provides both the precision afforded by thorough enumeration and the computational tractability of sample-based generation, all while leveraging batched execution to prevent overwhelming system memory.
Moreover, in large-scale experiments involving Monte Carlo sampling, storing all candidate randomizations in memory quickly becomes infeasible as $n$ and the total number of candidate randomizations grow. To address this, our implementation employs a streamlined approach where each batch of candidate randomizations is generated and evaluated, with only key information being preserved for the final evaluation step. This keeps the memory footprint low. The result is a scalable procedure that supports large experiments, since we avoid loading or caching an unwieldy number of candidate assignments in a single data structure. This key-generation process is visualized in Figure 3.
Figure 3: Visualization of key usage to avoid memory overhead associated with storing full treatment assignment vectors.
A Brief Model of Memory Gains from Key-Only Storage
To illustrate the gains from this key-only based approach, we here sketch a simple model. Let $n$ denote the number of units in the experiment, and suppose we wish to generate $M$ candidate randomization vectors $\{W^{(1)},\ldots,W^{(M)}\}$, where each $W^{(m)}$ is an $n$-dimensional binary treatment assignment vector (e.g., indicating which units are treated). Storing all $M$ generated vectors in memory requires space proportional to
$\underbrace{M\times n}_{\text{Total length of stored randomizations}}.$
Instead of storing every candidate randomization vector directly, one may store only an associated key $\kappa^{(m)}$ for each $W^{(m)}$. Here, $\kappa^{(m)}\in\{\mathbb{N}_{0}\}^{L}$ is a short seed (or split key) that can be used by a pseudorandom number generator (PRNG) to regenerate the corresponding vector $W^{(m)}$ on demand:
$W^{(m)}\;=\;r(\mathrm{PRNG}(\kappa^{(m)})),$
where $r(\cdot)$ is a function returning the full randomization from the key input. The key length, $L$, is much smaller than $n$ (usually, $L=2$). Hence, storing all $M$ keys occupies only
$\underbrace{M\times L}_{\text{Total length of stored keys}}$
memory units, which can be orders of magnitude smaller than $M\times n$. This sets up a memory reduction proportional to:
[…]
For an experiment with 1000 units, this ratio would be $1000/2$, indicating memory usage improvements of 500 times. At any later point, to reconstruct the accepted $W^{(m)}$ values for final analysis, we simply apply the PRNG again to these stored keys.
Hence, the key-only strategy offers several gains. First, we need not hold all vectors $W^{(m)}$ in memory; each $W^{(m)}$ can be generated briefly, evaluated, and discarded, while its key $\kappa^{(m)}$ is stored for possible use. This drastically reduces persistent storage. In large experiments (large $n$, large $M$), the overall memory requirement can thus decrease by a factor of $\frac{n}{L}$, enabling more efficient handling of massive randomization sets without sacrificing the ability to reproduce and analyze the accepted designs.
We next cover additional computational aspects supporting fastrerandomize.
3.2 Approximations
When evaluating multivariate covariate balance, the package provides a default threshold function based on Hotelling’s $T^{2}$. For very large matrices, an approximate inverse of the sample covariance matrix can be used (approximate_inv = TRUE), reducing computational overhead from matrix inversion while still enabling balance checking.
3.3 Just-in-Time Compilation with Precision Management
Each internal function (such as FastHotel2T2 and FastDiffInMeans) is decorated with JAX just-in-time (JIT) compilation, which translates high-level operations into optimized lower-level kernels for better performance of batch sampling (i.e., where the same computation is performed repeatedly). In addition, the package leverages JAX’s autovectorization (e.g., vmap) to automatically parallelize computations across candidate randomizations, further improving efficiency in large-scale designs. To further reduce memory use, computations can take advantage of low-precision data types, such as 16-bit floats for operations like matrix multiplication or aggregation operations, which can help conserve memory and accelerate numerical operations without significantly impacting the rerandomization criteria.
3.4 Hardware Acceleration
To further address the computational demands of rerandomization involving many units or covariates, fastrerandomize leverages GPU acceleration to mitigate runtime bottlenecks during both randomization generation and randomization-based inference. GPUs are particularly well-suited for parallelizing the small, repeated tasks that characterize rerandomization, such as repeatedly sampling and evaluating balance statistics for large pools of candidate randomizations.
Figure 4: High-level schematic of hardware acceleration of large-scale rerandomization. The CPU batches and sends candidate assignments to the GPU for massively parallel balance checks (e.g., Hotelling’s $T^{2}$). The GPU returns computed statistics for acceptance or rejection, leading to a final pool of acceptable assignments.
The package’s GPU acceleration strategy centers on efficient parallel processing of balance checks. As illustrated in Figure 4, the CPU first enumerates or samples assignment vectors in batches before offloading them to the GPU for parallel evaluation. The GPU then computes balance statistics (such as Hotelling’s $T^{2}$) across many candidate assignments simultaneously before returning results to the CPU, which accepts or rejects assignments based on the specified covariate-balance threshold. This parallel processing becomes especially advantageous as either the dimension of the covariate space or the number of Monte Carlo draws increases, since matrix operations can be performed concurrently across potentially tens of thousands of GPU cores.
Figure 5: Comparing hardware-accelerated performance of rerandomization steps on the CPU versus GPU. Bars show elapsed times (seconds) for (top) generating large pools of randomizations and (bottom) conducting the randomization-based inference. Left/right varies the maximum number of Monte Carlo draws (either $1\times 10^{5}$ or $2\times 10^{5}$). The $x$-axis indicates the number of covariates, the $y$-axis runtime. GPU acceleration provides substantial speed-ups in both the generation and testing phases, particularly as the number of covariates or draws increases. Here, $n$=1000; the CPU backend is an Apple M series chip (14 core); the GPU is 40-core METAL backend device.
Figure 5 illustrates how GPU acceleration yields substantial performance gains relative to the CPU for both randomization generation and randomization testing. With a larger number of covariates ($1000$) or more intensive Monte Carlo sampling ($2\times 10^{5}$ draws), GPU acceleration reduces the elapsed time dramatically by executing linear algebra and sampling tasks across thousands of cores in parallel. This allows the software to handle large-scale rerandomization designs that would otherwise be computationally prohibitive under a CPU-only workflow. (Note, though, that CPU-only users can still benefit from JIT-compilation and automatic vectorization capabilities in the package.)
Next, Figure 6 illustrates how the GPU-vs.-CPU speedup (i.e., how many times faster GPU is relative to CPU) increases when moving from $n=100$ to $n=1000$. On the horizontal axis, we vary the number of covariates; on the vertical axis, we plot the ratio of GPU speedups:
$\frac{\bigl{(}\text{CPU time}/\text{GPU time}\bigr{)}_{\!n=1000}}{\bigl{(}\text{CPU time}/\text{GPU time}\bigr{)}_{\!n=100}}.$
Values above 1 indicate the GPU’s performance advantage becomes larger at $n=1000$ than at $n=100$. The figure is faceted by the two main package steps — generate_randomizations vs. randomization_test—and by the number of Monte Carlo draws ($1\times 10^{5}$ or $2\times 10^{5}$). As the problem size grows, the ratio typically exceeds 1 (particularly so for the more computationally expensive candidate randomization generation step), indicating that the GPU’s relative speedup improves with larger $n$.
Figure 6: Comparison of GPU speedups at $n=1000$ vs. $n=100$. Bars show the ratio $\bigl{[}\text{GPU speedup at }n=1000\bigr{]}/\bigl{[}\text{GPU speedup at }n=100\bigr{]}$, faceted by the number of Monte Carlo draws ($1\times 10^{5}$ vs. $2\times 10^{5}$) and by the two main steps (generate_randomizations and randomization_test). A value above 1 indicates that the GPU’s advantage is larger at $n=1000$ than at $n=100$.
In the next section, we show how these speedups translate to practical usage within fastrerandomize, detailing its installation, core functions, and recommended workflows for both small- and large-scale experiments.
4 FastRerandomize Package Use
fastrerandomize provides an interface for researchers to (1) install and configure a hardware-accelerated JAX backend, (2) generate pools of acceptable randomizations via exact enumeration or Monte Carlo sampling, and (3) perform randomization-based inference respecting rerandomization. This section outlines the basic steps to install and initialize the software, generate randomizations, and conduct inference, followed by advanced usage tips for large-scale designs and pre-analysis evaluation.
4.1 Package Installation and Loading
Installing fastrerandomize from GitHub requires the devtools package:
# Install devtools if needed
# install.packages("devtools")
# Install fastrerandomize
devtools::install_github(
repo = "cjerzak/fastrerandomize-software/fastrerandomize"
# Load the package
library(fastrerandomize)
The package automatically calls out to JAX for hardware-accelerated computations. By default, fastrerandomize attempts to locate an existing conda environment named "fastrerandomize". To create or update this environment with all necessary backend packages, run:
build_backend(
conda_env = "fastrerandomize" # Or a name of your choice
This function installs numpy, jax, and other dependencies in a minimal Python 3.11 environment via conda. If you have a CUDA-capable GPU and drivers properly configured, jaxlib with CUDA support can be employed. On Mac M+ architectures, jax-metal is installed for METAL-based GPU acceleration.
4.2 Key Functions Overview
Below is an overview of the core fastrerandomize functions and their primary purposes:
build_backend()
Creates or updates a conda environment for running the hardware-accelerated JAX backend. By default, the environment is named "fastrerandomize". Internally, this function installs (or upgrades) jax, jaxlib, and related dependencies in a lightweight Python 3.11 environment. On machines with NVIDIA GPUs, a CUDA-enabled jaxlib is installed automatically if CUDA drivers are detected; on Apple Silicon (M1/M2), the METAL-accelerated jax is installed instead. This function is typically run once per session before calling any other fastrerandomize routines. The key arguments include:
  • conda_env: Name of the conda environment to create or update. Defaults to "fastrerandomize".
  • python_version: Optionally specify a Python version (default is "3.11").
  • force: Logical. If TRUE, forces an environment rebuild even if one already exists.
Example usage:
build_backend(conda_env = "fastrerandomize")
generate_randomizations()
Produces a pool of acceptable randomizations (treatment assignments) under a specified experimental design, employing either exact enumeration or Monte Carlo sampling. Internally, the function leverages JAX-based just-in-time (JIT) compilation to accelerate repeated operations (like random sampling and covariate-balance checks) on CPU or GPU. Accepted designs are retained based on a user-specified rerandomization threshold (e.g., Hotelling’s $T^{2}$ or a custom balance statistic). The output is returned as an S3 object of class fastrerandomize_randomizations, which has its own summary() and plot() methods.
Key arguments include:
  • n_units, n_treated: Integer counts specifying total units and the number assigned to treatment, respectively.
  • X: A numeric matrix of covariates (rows indicate units, columns indicate covariates).
  • randomization_type: Either "exact" (for smaller-scale experiments) or "monte_carlo" (for larger or high-dimensional designs).
  • randomization_accept_prob: Probability (between 0 and 1) controlling which fraction of candidate randomizations are accepted based on balance. Essentially, randomization_accept_prob = 0.01 means we keep the 1% best-balanced draws.
  • max_draws: Maximum total randomizations to generate (Monte Carlo only).
  • batch_size: Number of candidate randomizations generated per batch; controls memory usage.
  • approximate_inv: If TRUE, uses an approximate inverse for high-dimensional covariate balance calculations (e.g., large covariance matrix).
  • file: Optional filename to which accepted randomizations are written in CSV format; helps preserve memory by not returning them to R.
Example usage:
rand_mc <- generate_randomizations(
n_units = nrow(big_covariate_matrix),
n_treated = nrow(big_covariate_matrix)/2,
X = big_covariate_matrix,
randomization_type = "monte_carlo",
randomization_accept_prob= 0.01,
max_draws = 5e5,
batch_size = 1e4
The returned object, rand_mc, contains:
  • randomizations: An r by $n$ matrix of acceptable assignment vectors.
  • balance: The corresponding balance measure for each accepted randomization.
  • covariates, design_info, and other metadata.
S3 generics summary() and plot() let you inspect the balance distributions or the acceptance rates.
randomization_test()
Implements randomization-based inference for a given observed assignment and outcome vector, respecting the subset of randomizations accepted by rerandomization. This yields valid p-values and optionally fiducial intervals for the treatment effect under a sharp null hypothesis (e.g., $\tau=0$). Internally, it permutes outcomes across the accepted set of treatment assignments and compares the observed test statistic to its randomization distribution.
Key arguments include:
  • obsW: The observed assignment vector of length $n$ (0/1 for control/treatment).
  • obsY: Observed outcomes of length $n$.
  • candidate_randomizations: A matrix of dimension $r\times n$, containing the accepted randomizations generated by generate_randomizations().
  • findFI: Logical. If TRUE, attempts to compute a rerandomization-based fiducial interval (through test inversion).
  • test_statistic: Function or string specifying the test statistic; defaults to the difference-in-means.
Example usage:
test_results <- randomization_test(
obsW = rand_mc\$randomizations[1, ],
obsY = my_outcomes,
candidate_randomizations = rand_mc\$randomizations,
findFI = TRUE
The returned object has components:
  • p_value: A valid p-value for the test of no treatment effect, conditioned on rerandomization.
  • tau_obs: The observed difference-in-means (or other user-specified statistic).
  • FI: A fiducial interval (NULL if findFI = FALSE).
  • stat_distribution: The distribution of test statistics from the permuted outcomes.
S3 generics summary() and plot() provide summary and diagnostic plots (e.g., the randomization distribution with the observed statistic).
summary() and plot() S3 Methods
The S3 methods summary(), print(), and plot() are provided for objects returned by both generate_randomizations() and randomization_test(). They facilitate quick evaluation of the results and help diagnose potential issues (e.g., if not enough randomizations are retained or if balance thresholds are overly strict).
Typical usage:
# After generating randomizations
summary(rand_mc)
print(rand_mc)
plot(rand_mc)
# After randomization inference
summary(test_results)
print(test_results)
plot(test_results)
For fastrerandomize_randomizations objects, summary() reports basic statistics about the accepted pool (e.g., number of randomizations, summary of balance measures), while plot() visualizes the distribution of covariate-balance statistics. For fastrerandomize_test objects, summary() provides the p-value, observed effect, and (optionally) the fiducial interval, whereas plot() visualizes the permutation-based distribution of the test statistic.
Together, these functions support an end-to-end workflow for rerandomization:
  1. 1.
    Building a JAX-accelerated backend with build_backend();
  2. 2.
    Generating randomization pools with generate_randomizations();
  3. 3.
    Performing inference that respects rerandomization constraints via randomization_test();
  4. 4.
    Summarizing and visualizing each step with the accompanying S3 summary() and plot() methods.
Below, we illustrate these usage patterns in practice before later presenting two substantive applications involving real RCT data.
4.3 Workflow
1. Building the JAX backend.
When first working with the fastrerandomize packages, users should construct the environment and install JAX using build_backend():
build_backend(conda_env = "fastrerandomize")
2. Loading data.
After building the backend, prepare your data. We will use simulated data here for simplicity:
# Suppose we have n=20 units, with 10 to be treated
n_units <- 20
n_treated <- 10
# Example covariates: 5 features drawn from a standard normal
set.seed(12345)
X <- matrix(rnorm(n_units*5), nrow = n_units)
3. Generating randomizations.
Use generate_randomizations() to enumerate or sample acceptable assignments. By default, the function uses Monte Carlo sampling:
# Monte Carlo approach: draws random assignments and filters by covariate balance
rand_mc <- generate_randomizations(
n_units = n_units,
n_treated = n_treated,
X = X,
randomization_accept_prob= 0.01, # Keep top 1% most balanced
max_draws = 1e5, # Up to 100k draws
batch_size = 1e4, # 10k at a time
randomization_type = "monte_carlo",
verbose = TRUE
# calling S3 functions on the randomizations object
print( rand_mc )
# Number of candidate randomizations: 1000
# Number of units (columns): 20
summary( rand_mc )
# Balance statistics (summary):
# Min. 1st Qu. Median Mean 3rd Qu. Max.
# 0.06317 0.37250 0.48474 0.45728 0.57373 0.63574
plot( rand_mc )
Figure 7: Result of plot(rand_mc).
For small experiments (e.g., if $\binom{n}{n_{\textrm{Treated}}}\leq 10^{6}$), exact enumeration may be feasible:
rand_exact <- generate_randomizations(
n_units = 10,
n_treated = 5,
X = matrix(rnorm(10*3), nrow=10),
randomization_accept_prob = 0.2,
randomization_type = "exact"
4. Conducting randomization inference.
Once we have a set of acceptable assignments, we can conduct randomization-based inference for a hypothesized treatment effect. Suppose our observed outcomes are generated by:
# True effect of 1
tau_true <- 1
# Generate an observed assignment from the pool:
obsW <- rand_mc\$randomizations[1,]
# Construct outcomes: Y_i = X_i * beta + tau_true*obsW + noise
beta <- rnorm(ncol(X))
obsY <-as.matrix(X) %*beta + obsW * tau_true + rnorm(n_units, sd = 0.5)
obsY <- as.numeric(obsY)
With these data, we perform a randomization test:
test_results <- randomization_test(
obsW = obsW,
obsY = obsY,
candidate_randomizations = rand_mc\$randomizations,
findFI = TRUE # request a fiducial interval
The returned object from randomization_test includes:
  • p_value: The $p$-value for testing the null hypothesis that $\tau=0$, conditioned on the rerandomization design.
  • tau_obs: The observed difference-in-means estimate.
  • FI: (Optional) A fiducial interval for the treatment effect.
# Summary
print(test_results)
# P-value: 0.039
# Observed effect (tau_obs): 2.206593
# Fiducial interval: [0.5870508, 3.746144]
# Summary
summary(test_results)
#P-value:
# 0.039
#Observed effect (tau_obs):
# 2.206593
#Fiducial Interval:
# [0.5870508, 3.746144]
# Visualize results (including the fiducial interval)
plot(test_results)
Figure 8: Result of plot(test_results).
4.4 Advanced Usage and Practical Considerations
Selecting acceptance thresholds.
Selecting an acceptance threshold,
  • randomization_accept_prob
balances tighter covariate balance (lower acceptance probability) against the size of the retained randomization pool. If the threshold is too restrictive, you might have too few randomizations for stable inference. Conversely, if it is too lenient, the statistical benefits of rerandomization are diminished. A recommended approach is to explore multiple acceptance probabilities and compare how each influences downstream inference.
Large-scale designs.
For experiments with thousands of units or more:
  • Use randomization_type = "monte_carlo" to avoid enumerating all randomizations.
  • Increase max_draws as needed, but consider the trade-off between computing time and the desired pool size.
  • Adjust batch_size to manage memory. Smaller batches reduce peak memory usage at the cost of somewhat slower runtime.
  • Use hardware acceleration when possible, as matrix operations can be substantially faster on GPUs.
Storage and file I/O.
When generating very large randomization pools, specifying the file parameter in generate_randomizations() can save results directly to a CSV without returning them as an R object:
generate_randomizations(
n_units = nrow(big_covariate_matrix),
n_treated = round(nrow(big_covariate_matrix)/2),
X = big_covariate_matrix,
randomization_accept_prob = 0.01,
max_draws = 5e6,
batch_size = 5e3,
file = "my_accepted_randomizations.csv"
You can then load only subsets of the file into R for downstream analysis, preserving memory. Similarly, the return_type parameter can be set to "jax" if your workflow stays mostly in JAX (advanced usage).
Balance measures.
By default, the package uses Hotelling’s $T^{2}$ as the rerandomization criterion. You can supply a different function to threshold_func in generate_randomizations() if you prefer alternative balance statistics. The function must be vectorized in JAX (via vmap), returning a numeric measure for each randomization.
Fiducial intervals.
When findFI = TRUE in randomization_test(), the package attempts to compute an exact rerandomization-based interval. For large sample sizes or extremely large randomization pools, this calculation can be time-consuming. It is often helpful to set a smaller acceptance threshold or restrict the size of the randomization set when first experimenting with fiducial interval estimation.
4.5 Advanced Use: Optimal Threshold Exploration
To systematically explore how different acceptance thresholds might impact inference, you can iterate over a grid of randomization_accept_prob values and generate randomizations for each. Then, evaluate the potential gains in terms of smaller $p$-values or narrower fiducial intervals. For instance, the snippet below simulates a small experiment repeatedly:
# Setup parameters
set.seed( [^9] L)
X <- matrix(rnorm(100*3), ncol = 3)
beta <- rnorm(ncol(X))
ep <- rnorm(nrow(X), sd = 0.5)
# Investigate role of acceptance probability
probs <- seq(0.001, 0.05, length.out = 5)
p_values <- numeric(length(probs))
# loop over probs
for(i in seq_along(probs)) {
rand_pool <- generate_randomizations(
n_units = nrow(X),
n_treated = nrow(X)/2,
X = X,
randomization_accept_prob = probs[i],
max_draws = 1e5,
batch_size = 1e4
# Generate synthetic obsW, obsY:
obsW <- rand_pool\$randomizations[1, ]
tau_true <- 1
obsY <- as.matrix(X) %*beta + obsW * tau_true + ep
obsY <- as.numeric(obsY)
# Randomization test
test_out <- randomization_test(
obsW = obsW,
obsY = obsY,
candidate_randomizations = rand_pool\$randomizations
p_values[i] <- test_out\$p_value
plot(probs, p_values, type="b",
main="Effect of acceptance threshold on p-values")
Figure 9: Pre-analysis optimal threshold analysis.
This pre-analysis evaluation (results visualized in Figure 9) can guide practical design decisions before data collection, helping researchers choose rerandomization thresholds that balance computational feasibility with improved covariate balance.
4.6 Summary of Workflow
Figure 1 and the examples above illustrate the recommended workflow:
  1. 1.
    Initialize and build the Python/JAX backend, enabling JIT compilation and optional TPU/GPU acceleration.
  2. 2.
    Simulate or load covariate data for the planned experiment.
  3. 3.
    Generate randomizations by exact enumeration (small $n$) or Monte Carlo sampling (large $n$), using compiled code and, if available, GPU-accelerated batching.
  4. 4.
    Collect the resulting set of acceptable assignments, which can be returned to R or saved to disk (again exploiting compiled JAX routines).
  5. 5.
    Perform randomization-based inference using randomization_test() for your observed outcomes, respecting the rerandomization design and reusing the same GPU-accelerated kernels.
By following these steps, researchers can incorporate rerandomization into their experimental workflow while harnessing hardware-accelerated computations to handle designs that would otherwise be prohibitively expensive computationally.
5 Examples
5.1 Substantive Example 1: An Agricultural Field Experiment
In what follows, we illustrate a complete workflow for an agricultural field experiment drawn from
  • Burchardi KB, Gulesci S, Lerva B, Suárez Serrato JC (2019). “Moral hazard: Experimental evidence from tenancy contracts.” The Quarterly Journal of Economics, 134(1), 281–347.
Suppose we have $\mathit{n}$ plots of land, each plot having various measured covariates associated with the tenants (i.e., number of children, household size, and so forth). We want to assign half of these plots to treatment while ensuring good covariate balance, and then conduct a randomization-based inference respecting that rerandomization.
# Setup: install, load package, initialize JAX
# If not already installed:
# install.packages("devtools")
# devtools::install_github("cjerzak/fastrerandomize-software/fastrerandomize")
library(fastrerandomize)
# Optional: build the Python/JAX backend in a conda environment
#build_backend(conda_env = "fastrerandomize")
# Obtain pre-treatment covariates
data(QJEData, package = "fastrerandomize")
myCovariates <- c("children","married","hh_size","hh_sexrat")
QJEData <- QJEData[!is.na(rowSums(QJEData[,myCovariates])),]
X <- QJEData[,myCovariates]
# Analysis parameters
n_units <- nrow(X)
n_treated <- nrow(X)/2
# Generating rerandomizations with a balance criterion
# Monte Carlo approach (better for large n)
rand_mc <- generate_randomizations(
n_units = n_units,
n_treated = n_treated,
X = X,
randomization_type = "monte_carlo",
randomization_accept_prob= 0.02, # keep top 2%
max_draws = 200000,
batch_size = 2000,
verbose = TRUE
print(rand_mc)
summary(rand_mc)
plot(rand_mc)
# Simulating observed outcomes
# Select the first accepted randomization from the Monte Carlo set
obsW <- rand_mc\$randomizations[1, ]
# True effect of 1, plus some linear dependence on X
tau_true <- 1
beta <- rnorm(ncol(X))
# Generate outcomes: Y = X * beta + W * tau + Normal(0, 0.5^2)
obsY <- as.matrix(X) %*beta + obsW * tau_true + rnorm(n_units, sd = 0.5)
obsY <- as.numeric(obsY)
# Randomization test for the observed assignment
test_out <- randomization_test(
obsW = obsW,
obsY = obsY,
candidate_randomizations = rand_mc\$randomizations,
findFI = TRUE # request a fiducial interval
print(test_out)
summary(test_out)
plot(test_out)
In this example, we generated synthetic data and then used two different methods ("exact" and "monte_carlo") for creating balanced treatment assignments. The subsequent randomization test respects the fact that only certain assignments were accepted due to covariate-balance constraints, ensuring valid inference. The p_value from randomization_test() and the optional fiducial interval (set findFI=TRUE) provide exact inferences under the specified rerandomization design.
Researchers working with more extensive field data (e.g., thousands of plots and covariates) can exploit GPU acceleration for both generating balanced assignment pools and running large-scale randomization tests. By adjusting parameters such as
  • randomization_accept_prob
and
  • max_draws
users can navigate the trade-off between computational feasibility and tighter covariate balance in larger experiments ($n$>20).
5.2 Substantive Example 2: An Anti-Poverty RCT
In the following example, we will show the benefits of using the package for randomized control trials where the number of covariates is large. In particular, we will draw on an anti-poverty RCT conducted in Uganda:
  • Blattman, C., Fiala, N. and Martinez, S. (2020). The Long-term Impacts of Grants on Poverty: Nine-year Evidence from Uganda’s Youth Opportunities Program. American Economic Review: Insights, 2(3), 287-304.
To enrich the original anti-poverty dataset, we followed the approach of Jerzak et al. (2023) and appended pre-treatment image features derived from satellite imagery centered on each RCT unit’s approximate geo-location. This augmentation yields over 500 unique covariates per unit from a fine-tuned CLIP image model (Lu et al., 2017), capturing a fine-grained depiction of pre-treatment contextual conditions (e.g., related to the environment or agriculture). We then assess the randomness of various treatment assignments in this high-dimensional space by computing the Mahalanobis distance, which serves as a robust measure of covariate balance across treatment and control groups.
The below codeblock will compare Mahalanobis distance values from the original experiment with what we would obtain via rerandomization in image space. A greater Mahalanobis distance value is an indicator that the covariate means for the treatment and control groups are very different, meaning that the randomization did not produce groups with well-balanced covariates. Thus, “good” randomizations ought to minimize their M-distance values. We will compare the M-distance value under the implemented and a hypothetical rerandomization-based design.
# Acquire package data for the Youth Opportunities Experiment
data( YOPData, package = "fastrerandomize")
#Make sure row dimensions match
dim(YOPData\$ImageEmbeddings)
dim(YOPData\$RCTData)
# For this tutorial, we assume you already have fastrerandomize installed:
library(fastrerandomize)
set.seed(543210)
# Load the built-in data, which has:
# 1) YOPData\$RCTData (contains treatment indicator + possibly other columns)
# 2) YOPData\$ImageEmbeddings (CLIP-based satellite imagery embeddings)
data(YOPData)
# Verify the dimensions
dim(YOPData\$ImageEmbeddings) # e.g., N x K matrix of embeddings
dim(YOPData\$RCTData) # e.g., N x a few columns (including "treated")
# 1) Define a function to compute Mahalanobis distance
# between average treated and control covariates
compute_m_dist <- function(merged_assignment) {
library(dplyr)
# merged_assignment is a data.frame or matrix whose first column is "treated"
# and the remaining columns are the covariates/embeddings.
# The function then computes the sample mean difference,
# the pooled covariance matrix, and uses mahalanobis() to get the M-dist.
treated <- merged_assignment %>filter(treated == 1) %>select(-treated)
control <- merged_assignment %>filter(treated == 0) %>select(-treated)
m_distance <- mahalanobis(center = colMeans(treated),
x = colMeans(control),
cov = diag(diag(cov(rbind(treated, control)))))
return(m_distance)
# 2) Generate a rerandomization-based design:
# We’ll accept random assignments to T=1 or T=0
# only if they pass a certain (low) acceptance probability.
# This is just a demonstration: randomization_accept_prob = 0.0001
candidate_randomizations_array <- generate_randomizations(
n_units = length(YOPData\$RCTData\$treated),
n_treated = sum(YOPData\$RCTData\$treated),
X = YOPData\$ImageEmbeddings,
randomization_accept_prob = 0.0001, # strict acceptance
max_draws = 10^5,
batch_size = 10^4,
approximate_inv = TRUE
# This returns an S3 object of class "fastrerandomize_randomizations",
# which has (1) randomizations, (2) balance, etc.
cat("Number of accepted randomizations:",
[…]
# Compare M-distances
# (A) M-distance of a purely random assignment
# 1) random draw of 0/1
# 2) cbind to embeddings
# 3) compute M-distance
random_array <- sample(YOPData\$RCTData\$treated,
size = nrow(YOPData\$ImageEmbeddings),
replace = FALSE)
merged_arr <- cbind("treated" = random_array,
YOPData\$ImageEmbeddings)
random_m_distance <- compute_m_dist(as.data.frame(merged_arr))
# (B) M-distance of the YOP Data’s actual assignment
# We assume the column in RCTData is called "treated" (if not, rename below).
merged_assignment <- cbind("treated" = YOPData\$RCTData\$treated,
YOPData\$ImageEmbeddings)
given_m_distance <- compute_m_dist(as.data.frame(merged_assignment))
# (C) Average M-distance across the accepted randomizations
# We’ll loop over some of the accepted
# assignment vector in candidate_randomizations_array
# and compute the M-dist. Then we compute the average
nM <- 100 # how many accepted to analyze
SampIndices <- sample(1:nrow(candidate_randomizations_array\$randomizations),
nM,
replace = T)
mean_m_dist <- 0
mDist_vec <- c()
for (i in SampIndices){
# For each row i, we have a length-N vector of 0/1 for the assignment
merged_temp <- as.data.frame(
cbind("treated" = candidate_randomizations_array\$randomizations[i, ],
YOPData\$ImageEmbeddings))
mDist_ <- compute_m_dist(merged_temp)
mDist_vec <- c(mDist_vec, mDist_)
mean_m_dist <- mean_m_dist + mDist_ / nM
# print results
[…]
# Random assignment M-distance: 1.506205
[…]
# Actual YOPData assignment M-distance: 13.5678
[…]
# M-distance over accepted randomizations: 0.3384692
In terms of results, we see from these M-distance calculations that the purely random assignment yields a moderate degree of imbalance in the high-dimensional covariate space, while the actual YOP assignment exhibits an even larger imbalance (M-distance of 13.57). In contrast, the rerandomized assignments accepted by fastrerandomize achieve substantially lower M-distances (mean of 0.34 among the sampled accepted assignments).
We also found that the calculation of just $10^{2}$ randomization M distances in this high-dimensional case in (admittedly un-optimized) base R has a longer runtime than the hardware accelerated calculation of $10^{5}$ distances using fastrerandomize. These results highlight how hardware-accelerated rerandomization can markedly improve covariate balance when dealing with high-dimensional embeddings (e.g., hundreds of satellite-derived features). Simply put, the package can help unlock more stringent rerandomization designs for more complex, high-dimensional settings.
# Performing Randomization Tests
# Compute p-value of assuming a rerandomization design random assignment
random_test_one <- fastrerandomize::randomization_test(
X = YOPData\$ImageEmbeddings,
obsY = YOPData\$RCTData[,"outcome"],
obsW = candidate_randomizations_array\$randomizations[1,],
candidate_randomizations = candidate_randomizations_array\$randomizations
random_test_one\$p_value # 0.2
summary(random_test_one)
Finally, we here see that the call to randomization_test() would produce a $p$-value of 0.20 for the sharp null of zero treatment effect under the assumption of a rerandomization-based design. This illustrates how fastrerandomize can integrate design-based inference with stringent acceptance of well-balanced random assignments in high-dimensional spaces.
6 Future Development Plans
Although the current iteration of fastrerandomize provides a robust and scalable tool for conducting rerandomization in large-scale experiments, several avenues for future development remain. Our primary goals include:
Further Performance Optimizations and Multi-Node Scaling. While the package currently benefits from single-device hardware acceleration, we plan to extend support to multi-GPU systems and distributed computing frameworks. Approaches such as message-passing interfaces could further facilitate parallelization across large, multi-node clusters, allowing users to handle experiments at an even greater scale. Integrating these frameworks is especially relevant for high-performance computing environments where tens or hundreds of GPUs may be available.
Extensions to Complex and Sequential Designs. Many real-world experiments go beyond a single wave of assignment, involving multi-armed, factorial, or sequentially adaptive designs (Robbins, 1952). A natural step is to incorporate rerandomization at multiple stages (e.g., mid-experiment reassignments or multi-armed bandit designs) while preserving valid randomization-based inference. Future versions of fastrerandomize will explore iterative balancing strategies and acceptance criteria for each stage of a multi-period or multi-armed experiment, potentially incorporating Bayesian or machine-learning-based methods for dynamic updating.
Quantum Speedups. Recent advances in quantum computing open up new possibilities for randomized algorithms and high-dimensional sampling (Google Quantum AI, 2024). Future research may investigate leveraging quantum computing for generating or filtering candidate assignments, especially in combinatorial tasks where quantum speedups would likely prove beneficial based on established results in amplitude amplification (Brassard et al., 2002). While these methods are not yet broadly accessible, experimental prototypes could showcase how quantum-enabled sampling routines scale with problem size and dimensionality, complementing the classical hardware acceleration already employed.
Integration with Other Experimental Design Packages. Finally, future releases will aim for tighter integration with existing design-focused R packages (e.g., DeclareDesign or randomizeR), enabling analysts to incorporate rerandomization more fluidly into broader workflows. Ideally, fastrerandomize would serve as a flexible backend for randomization generation and randomization-based inference, where other packages manage tasks such as power analysis, blocking, or multi-armed design specification.
By pursuing these directions, we aim to reinforce fastrerandomize as a comprehensive and extensible solution for large-scale experiments, addressing emerging methodological needs while offering next-generation computational performance.
7 Conclusion
The fastrerandomize package provides fast rerandomization and randomization inference by combining a statistically principled framework with hardware-accelerated computations. While the current release already supports large-scale designs by leveraging efficient memory operations and JAX backends (CPU, CUDA, and METAL), future work will focus on scaling to even larger randomization spaces by refining memory handling, exploring approximate methods that reduce the computational footprint for massive experiments, accommodating additional balance criteria to reflect more nuanced research needs, and broadening hardware acceleration capabilities across new platforms. By pursuing these directions, we aim to solidify fastrerandomize as a robust solution for researchers facing demanding experimental designs that require rigorous randomization-based inference.
Computational Details
Results obtained using fastrerandomize 0.1 with JAX backend version 0.4.38. R itself and supporting packages are available from CRAN at https://CRAN.R-project.org/.
License
The fastrerandomize package is distributed under the GPL-3 license; users retain the freedom to run, study, modify, and redistribute the software, in alignment with open-source principles.
Acknowledgments
Authors listed in alphabetical order. We thank Kaz Sakamoto and the R community for helpful feedback and testing.